cbagwe8982:

Sally is 54 years old and her mother is 80, how many years ago was Sally's mother three times her age?

surjithayer:

lt x years ago Sally's mother's age ie three times Sally. 80-x=3(54-x) solve for x.

SmokeyBrown:

Sally's age is 26 less than her mother's age because 80 - 54 = 26. In other words, Sally is 26 years younger than her mother. We can say S = M - 26; this will always be true, even if we go back any number of years. We want to find the age at which Sally's mother, M, was 3 times Sally's age or 3S = M At this point in time, it would still be true that S = M - 26 The trick is to find values of S and M such that both statements are true: 3S = M AND S = M - 26 Given that S = M - 26, we can use substitution to replace "S" in the first equation with (M - 26) and then solve for M: 3(M - 26) = M 3M - 78 = M 2M = 78 M = ??? It would then be obvious how many years ago that was, compared to M's current age of 80

@surjithayer wrote:
lt x years ago Sally's mother's age ie three times Sally. 80-x=3(54-x) solve for x.
I think surjithayer's method also works, and it is a little more straightforward :)